### v1

# Harnessing quantum back-action for time-series processing

Abstract
Quantum measurements affect the state of the observed systems via back-action. While projective measurements extract maximal classical information, they drastically alter the system. In contrast, weak measurements balance information extraction with the degree of disturbance. Considering the prevalent use of projective measurements in quantum computing and communication protocols, the potential benefits of weak measurements in these fields remain largely unexplored. In this work, we demonstrate that incorporating weak measurements into a quantum machine-learning protocol known as quantum reservoir computing provides advantages in both execution time scaling and overall performance. We analyze different measurement settings by varying the measurement strength across two benchmarking tasks. Our results reveal that carefully optimizing both the reservoir Hamiltonian parameters and the measurement strength can significantly improve the quantum reservoir computing algorithm performance. This work provides a comprehensive and practical recipe to promote the implementation of weak measurement-based protocols in quantum reservoir computing. Moreover, our findings motivate further exploration of experimental protocols that leverage the back-action effects of weak measurements.
preprint:
I Introduction
A fundamental aspect that makes quantum measurements unique and counter-intuitive from a classical perspective is back-action, the influence of observers on the state of quantum systems. From a foundational perspective, this phenomenon is central to the ‘measurement problem’, sparking formal and philosophical debates about different interpretations of quantum mechanics since von Neumann’s time [von1955mathematical, wheeler2014quantum]. From a more applied viewpoint, the so-called collapse of the wave function after the measurement process is crucial for different quantum technologies. For instance, in quantum key distribution [BENNETT20147], the disturbance caused by measurement is exploited to detect eavesdropping attempts, ensuring secure communication. Similarly, quantum sensing [degen2017quantum] harnesses the extreme sensitivity of quantum systems to external interactions, achieving precision levels beyond classical methods. In this work, we explore a novel application of quantum back-action: enhancing the performance of machine-learning algorithms.
The degree of disturbance induced by a measurement depends on its nature. Conventional projective measurements cause substantial disruption to the quantum system, forcing it into the eigenstate (or a degenerate subspace) associated with the measured outcome. This strong interaction collapses superposition and fully determines the system’s state. In contrast, weak measurements provide a gentler alternative by using indirect interaction between the system and the measurement apparatus [10.1093/acprof:oso/9780199213900.001.0001]. The strength of the coupling and the duration of this interaction govern both the weakness of the measurement and the amount of information extracted. Consequently, weak measurements provide only partial information about the system, inducing a back-action that depends on the degree of measurement weakness. This approach allows the quantum state to retain some coherence, enabling ongoing observation of the system without fully collapsing it. This makes weak measurements a valuable tool for quantum information processing [PhysRevA.62.012307, SINGH2014141]. Weak measurements have been both demonstrated and widely used across various quantum platforms [pan2020weak, spagnolo2022experimental, murch2013observing]. Superconducting qubits, for example, are typically measured through weak measurements using ancillary microwave pulses [naghiloo2019introduction, weber2016quantum]. By adjusting the strength and duration of these pulses, experimenters can tune the measurement strength from weak to strong, with the latter approaching projective measurements in the limit [murch2013observing, Murch2016, doi:10.1126/science.1226897]. In general, the capacity to implement weak measurements relies on the ability to prepare convenient ancillas [Yasuda2023] or to extract information through projectively measuring only a portion of the system [hu2024overcoming, viqueira2023density].
Figure 1: Quantum reservoir computing via weak measurements. (a) Example of the reservoir’s prediction (dashed lines) over the target series (solid lines) on two specific instances ($\eta=5$) for the forward prediction task (top panel) and memory retrieval task (bottom panel) (see section II.2). (b) Schematic view of the information processing step of the QRC algorithm, the coupling strength $g$ between the measurement apparatus and the reservoir is the main parameter of this work’s study. (c) Illustrative example of the effect of weak measurement on a single spin state for different values of the measurement strength, $g$. As $g\to\infty$, the measurement becomes equivalent to a projective measurement, fully collapsing the state onto one of the measurement axes. However, as $g$ decreases, the post-measurement state is no longer collapsed along the measurement axis and is less perturbed. In this analysis, we assume measurements are performed along the $z$-axis.
In this work, we demonstrate that adjusting the strength of quantum measurements enables the optimization of quantum back-action introduced into the system’s evolution, thereby enhancing the performance of machine learning tasks beyond what is achievable without this control. We illustrate our findings within the context of time-series processing using quantum reservoir computing (QRC) [mujal2021opportunities]. QRC belongs to the broader category of quantum machine learning, a field that emerged from the desire to combine machine learning techniques with the unique advantages of quantum systems [dawid2023]. First introduced in [fujii2017harnessing], it is a supervised quantum machine learning technique that leverages the advantages of quantum systems as the exponential growth of their state space with increasing system size. At the same time, it incorporates the simple and efficient training strategy characteristic of classical reservoir computing [jaeger2004harnessing], where the internal dynamics of the reservoir are fixed and only the output layer is trained. Recently, there has been growing interest in QRC and related approaches, with progress both in theoretical advancements [Nakajima2019boosting, Ghosh2019, Ghosh2019prl, Kutvonen2020, martinez2020information, Ghosh2020reconstructing, Tran2020higherorder, Nokkala2021, martinez2021dynamical, Ghosh2021neuromorphic, mujal2021analytical, govia2022nonlinear, Angelatos2021qstatemeas, Nokkala2022, Llodra2022, Domingo2022, Yelin2022, Nokkala2023, GarciaBeni2023, martinez2023finitedim, Krisnanda2023tomographic, mujal2023time, Innocenti2023, Wudarski2023, GarciaBeni2024squeezing, sannia2024dissipation, Nokkala2024, Jaurigue2024memory, Li2024estimating, Kobayashi2024, Domingo2024, Kobayashi2024eco, palacios2024role, Llodra2024] and experimental implementations on diverse platforms [negoro2018machine, Chen2020, Suzuki2022, spagnolo2022experimental, Pfeffer2022, Kubota2023, Yasuda2023, hu2024overcoming, Paternostro2024, Yelin2024]
In QRC, memory refers to the reservoir’s ability to retain and utilize information from past inputs to inform future predictions. It plays a key role in processing temporal series by capturing the time correlations within the dataset. The amount of memory retained by quantum reservoirs has been shown to depend on their dynamical regime [martinez2021dynamical] and is reduced by the effects of quantum back-action [mujal2023time]. When online processing time series, projective measurements introduce significant decoherence, which negatively impacts the memory capacity of the reservoir. As a result, to truly take advantage of the scaling benefits in online processing, a weak measurement approach is more suitable [mujal2023time]. With weak measurements, there is a tradeoff between the amount of information extracted and the quantum back-action introduced. While this reduces disturbance to the reservoir’s memory, it also limits the information obtained, leaving the potential advantage uncertain. This approach has already attracted experimental interest and has been realized in superconducting devices through mid-circuit measurements [hu2024overcoming, Yasuda2023].
Interestingly, in various contexts, controlling the degree of decoherence introduced into the reservoir has been shown to improve its memory capacity [martinez2023finitedim, sannia2024dissipation, monzani2024leveraging, palacios2024role]. In this work, we build on this insight by applying it to a weak measurement-based protocol [mujal2023time]. By leveraging the ability to tune the measurement strength, we optimize the tradeoff between information extraction and back-action. This allows for an online processing scheme that benefits from linear scaling without suffering drawbacks from decoherence. Instead, by carefully controlling the measurement strength, we can enhance the reservoir’s capacity. To achieve this, we introduce a new figure of merit to compare the performance of weak measurement-based protocols, where back-action is controlled, with projective measurement-based protocols that rely on restarting strategies to avoid back-action. By optimizing this metric, we identify the optimal measurement strength that maximizes performance. We illustrate this improvement across different realizations of reservoirs operating in distinct dynamical regimes, validated through two benchmark tasks: the forecasting of the Santa Fe chaotic series [PhysRevA.40.6354] and the short-term memory task . Furthermore, we extend our analysis to a realistic finite resource scenario and show that the enhancement effect is amplified by the favourable scaling properties of the weak measurement protocol.
II Preliminaries
II.1 Quantum reservoir computing via weak measurements
Reservoir computing was proposed as a training-efficient alternative to more complicated recurrent neural network models like Long Short-Term Memory for processing temporal data and time series [jaeger2004harnessing]. Quantum reservoir computing extends the concept by leveraging quantum properties and dynamics, using a quantum system as the reservoir [fujii2017harnessing]. The general pipeline of QRC algorithms is structured in two steps: first the reservoir has to process the time series of interest and then a linear regression model is trained to perform predictions (Fig. 1 (a)). While the latter is a straightforward process, the former includes three stages to be repeated for each element of the dataset: input encoding, time evolution and information extraction (Fig. 1 (b)). At the encoding stage, the information of the $k$-th element of the time series is injected into the amplitudes of the state of a fixed spin of the reservoir [fujii2017harnessing, martinez2021dynamical, mujal2021analytical]. The information is then spread through the system by evolving it for a time $\Delta t$ under a transverse field Ising Hamiltonian $\hat{H}$. Ultimately, we measure a set of observables $\{\langle\hat{O}_{i}\rangle\}$ to extract information from the quantum system. However, when performing projective measurements, it is convenient to reset the system and reprocess the previous $k$ elements without applying measurements before processing the next input, $k+1$. This is because projective measurements erase the stored information in quantum coherences through a destructive process. This procedure, known as the restarting protocol (RSP), is the most common practice in the research community. Recently, a novel protocol was introduced that employs weak measurements instead of projective ones [mujal2023time]. By using weak measurements, only a partial amount of information is lost due to the back-action effect (Fig. 1 (c)), allowing the QRC algorithm to run continuously without restarting at each step in the time series processing. This approach is referred to as the online protocol (OLP) [mujal2023time]. The OLP introduces back-action during the evolution as a tradeoff, allowing for better algorithm scaling by avoiding the need to restart at each input.
II.2 Two benchmark studies
We perform an in-depth analysis of how the back-action effect due to using a weak measurement-based protocol affects the performance of QRC. To do so, we take into account two tasks: the forecasting of a chaotic time series and the efficiency in storing past information. The chaotic time series we use in this work is the Santa Fe laser time series, a common benchmark dataset that follows the chaotic behavior of a $NH_{3}$ laser [PhysRevA.40.6354]. On the other hand, to test the ability to retrieve past information through the short-term memory task, we generate a set of 1000 random points from the uniform distribution in the range $[0,1]$. We build 20 different sub-tasks for both benchmarks where $\eta$, the prediction distance, varies from $1$ to $20$. Given an $\eta$, the target series $k$-th element becomes $y_{k}(\eta)=s_{k\pm\eta}$, where $\vec{s}$ is the original time series. The capacity of the reservoir for each sub-task is evaluated by computing the coefficient of determination $C(\eta)$ of the algorithm predicted series $\vec{\mathtt{y}}$ and the sub-task corresponding true one $\vec{y}(\eta)$:
$C(\eta)=\frac{\text{cov}^{2}(\vec{\mathtt{y}},\vec{y}(\eta))}{\text{var}(\vec{\mathtt{y}})\text{var}(\vec{y}(\eta))}.$ (1)
The predictions of the QRC algorithm are generated by a linear model trained on $70\%$ of the output dataset, which consists of the expectation values of the measured observables for each time series element. An example of the predictions of the QRC algorithm on a specific instance of the two tasks is presented in Figure 1 (a). Then, the overall performance of the reservoir on the task is measured through the sum capacity:
$C_{\Sigma}=\sum_{\eta=1}^{20}C(\eta).$ (2)
III Methods
III.1 Reservoir Hamiltonian
The quantum reservoir considered in this work consists of a system of $N$ spins with an all-to-all connection scheme, where the dynamics is governed by the transverse-field Ising Hamiltonian:
$\hat{H}=\sum_{i<j=0}^{N-1}J_{ij}\hat{\sigma}^{x}_{i}\hat{\sigma}^{x}_{j}+\frac{h}{2}\sum_{i=0}^{N-1}\hat{\sigma}^{z}_{i}.$ (3)
In the numerical simulations, we considered systems of $N=6$ spins with random couplings $J_{ij}$ uniformly sorted from the range $[-J/2,J/2]$ and we set $J=1$. Then, when referring to the other parameters of the Hamiltonian we always implicitly express them in units of $J$. At the beginning of the protocol, the initial state of the system is set to be $|\Psi\rangle=|0\rangle^{\otimes N}$.
III.2 Details of the algorithm pipeline
Focusing on the first step of the QRC algorithm pipeline, we provide a detailed description of the framework used in this work. To process the time series, firstly is necessary to normalize it and encode each element in the quantum system. Among the $N$ spins of the reservoir, we consider the following bipartition: we fix one spin to act as the input node, partition $A$, while the others $N-1$ form the partition $B$. For each element of the sequence $s_{k}$, the encoding routine consists of resetting $\hat{\rho}^{A}$ to the pure state $|\psi_{k}\rangle=\sqrt{1-s_{k}}|0\rangle+\sqrt{s_{k}}|1\rangle$ [mujal2021analytical]. After the input injection, the reservoir evolves under $\hat{H}$ for a time $\Delta t=10/J$, $\hat{U}=e^{-i\hat{H}\Delta t}$. In summary, if before the encoding of the element $s_{k}$ the state of the reservoir was $\hat{\rho}_{k-1}$, afterward the reservoir undergoes a completely positive trace-preserving (CPTP) map [fujii2017harnessing, martinez2021dynamical]:
$\mathcal{L}_{k}(\hat{\rho}_{k-1})=\hat{U}(|\psi_{k}\rangle\!\langle\psi_{k}|\otimes{\rm Tr_{A}}(\hat{\rho}_{k-1}))\hat{U}^{\dagger}.$ (4)
At this point, before the next input injection, it is necessary to extract the elaborated information from the system. As previously described, here we use the online protocol proposed in [mujal2023time] based on weak measurements. The effect of the measurement on the $z$-axis direction is modeled through the action of the matrix $\hat{M}=(\mathbf{I}+e^{-\frac{g^{2}}{2}}\hat{\sigma}_{x})^{\otimes N}$, such that the state after a measurement with strength $g$ is given by:
$\hat{\rho}_{k}=\hat{M}\odot{\mathcal{L}}_{k}(\hat{\rho}_{k-1}),$ (5)
where $\odot$ indicates the element-wise product. Actually, the CPTP map in this last equation is equivalent to a dephasing channel [martinez2023finitedim]. We can then extract the expectation value of the observable $\hat{\sigma}_{i}^{z}$ on the i-th spin as ${\rm Tr}(\hat{\rho}_{k}\hat{\sigma}_{i}^{z})$. To measure single qubit observables in the $x$ and $y$ direction it is necessary to apply proper rotations to ${\mathcal{L}}_{k}(\hat{\rho}_{k-1})$ before applying $\hat{M}$. For two-spin observables the same description applies.
After having processed all the elements of the sequence $\vec{s}$, we obtain a dataset where for each element of the input data vector we have the corresponding expectation value of the single-spin observables $\langle\hat{\sigma}^{x}_{i}\rangle,\langle\hat{\sigma}^{y}_{i}\rangle,\langle\hat{\sigma}^{z}_{i}\rangle$ for each spin of the reservoir and the two-spin observables $\langle\hat{\sigma}^{x}_{i}\hat{\sigma}^{x}_{j}\rangle,\langle\hat{\sigma}^{y}_{i}\hat{\sigma}^{y}_{j}\rangle,\langle\hat{\sigma}^{z}_{i}\hat{\sigma}^{z}_{j}\rangle$ for each pair of spins of the reservoir. Since measuring in the $x$ and $y$ direction implies applying rotations, this means that the algorithm has to go through the entire dataset three times and not just once.
IV Results
IV.1 Performance analysis in the infinite resources scenario
Previous work [mujal2023time] revealed that the same performance achieved with the RSP can, in some cases, also be obtained with the OLP, despite the introduction of back-action effects. On the basis of this, we analyze the performance of QRC with the OLP at different measurement strengths. Moreover, we consider different external field strengths, $h$, i.e. different dynamical regimes of the reservoir [martinez2021dynamical]. Both quantities are expressed as dimensionless throughout this work.
To quantify the performance of one protocol with respect to the other, we introduce the performance ratio, $P_{R}$, which measures how the QRC algorithm running with the OLP compares with the analogous QRC algorithm with the RSP. Specifically, $P_{R}(g,h)$ is defined as the ratio between the OLP’s sum capacity, given a fixed $g$ and $h$, and the best sum capacity, $h$ wise, obtained with the RSP:
$P_{R}(g,h)=\frac{C_{\Sigma}^{\text{OLP}}(g,h)}{\text{max}_{h}(C_{\Sigma}^{\text{RSP}}(h))}$ (6)
We choose the best-performing RSP rather than the corresponding RSP performance for the same $h$, as we aim for a metric that indicates not only if the OLP performs as well as the RSP but also if the QRC algorithm performs well in general. While this is a strict metric, it avoids focusing on trivial scenarios where both protocols perform poorly. With this definition, a $P_{R}>1$ indicates a practical advantage of the OLP over the RSP. We average the $P_{R}$ over 50 different reservoir realizations, where the couplings in between the spins are randomly selected each time.
In Figure 2 we show how $P_{R}$ changes at different $g$ and $h$ values for the forward prediction task on the Santa Fe dataset (top panel), and for the short-term memory task (bottom panel). Each colormap is made up of 12 by 40 points and is computed by assuming that we have access to an infinite amount of resources, i.e. an infinite number of shots for each measured observable. The $P_{R}$ at $g=0$ is set to be 00 since it corresponds to the case in which the measurement apparatus is not interacting with the reservoir and the measurement output is then a random number.
Figure 2: Performance ratio analysis for the forward prediction task (top) and for the short-term memory task (bottom). A $P_{R}$ in the range $[0,1)$ shows that the best performing RSP is performing better than the OLP in that point of the $g-h$ map. A $P_{R}$ around 1 shows that the OLP is performing as good as the best performing RSP, and over 1 indicates an over-performance.
For both tasks, we can highlight the clear emergence of a region for which the OLP achieves comparable performances to the best-performing RSP. In the case of the forward prediction task, this region is spread along the $h$ axis while for the short-term memory task is peaked for $h$ close to $10^{-1}$. Moreover, by looking at the top panel we reveal a region in the parameter space where there is a numerical emergence of the OLP overperforming the RSP, i.e. a red region, for a $g$ value close to $1$. In this case, the back-action effect, which in principle is expected to negatively affect the performance of QRC, brings instead a positive contribution that allows the maximum performance of the RSP to be overcome.
Hence, the emergence of these white and red regions highlights that a weak measurement-based protocol, if properly tuned, loses nothing in terms of performance compared to the RSP. To stress this last point, in Figure 3, we compare the different sum capacities of the QRC running with the RSP and with the OLP in the case in which $g$ is fixed to a non-optimal value and in the case in which it is fixed to the value that maximized $P_{R}$ in Figure 2. As a first observation, we can refer to [martinez2021dynamical] to justify the dependence of the performance on the $h$ value. In fact, along the horizontal axis, the quantum system undergoes a transition from a localized regime to an ergodic one and it was showcased that being in the ergodic regime is a preferential condition to perform reservoir computing. In addition, what emerges clearly from these plots is that by maximizing the $P_{R}$ one can optimize the strength of the measurement such that the reservoir performance is not negatively affected by the back-action and making the OLP as reliable as the RSP in terms of performance.
Figure 3: Sum capacity comparison of the RSP with respect to the OLP with and without optimizing the strength of the measurement $g$. With $g^{*}$ we denote the $g$ value that maximizes the $P_{R}$. The top panel shows the results for the forward prediction task while the bottom panel is for the short-term memory task.
The previous analysis is related to the sum capacity of the QRC across two different tasks. However, for different sub-tasks, defined by varying $\eta$, the performance may systematically differ. In this context, Figure 4 compares two points in the $g-h$ parameter space explored in Fig. 2, the one that maximized $P_{R}$ and a non-optimal point. We then evaluate the performance for each different sub-task relative to the performance of the RSP at optimal $h$.
Figure 4: Capacity for different sub-tasks. The figure compares the performance in single sub-tasks of the optimal restarting protocol with respect to the optimal and non-optimal versions of the OLP. The top panel shows the results for the forward prediction task while the bottom panel is for the short-term memory task.
For the forward prediction task, we observe that, given an optimal external field and measurement strength, the OLP consistently outperforms the RSP for every value of $\eta$. In contrast, for the memory retrieval task, the optimal OLP outperforms the RSP only for $\eta$ values less than 10, while the RSP performs better for the other sub-tasks. Moreover, in the upper panel, a non-optimal choice of parameters results in a constant factor decrease in performance. Instead, for the short-term memory task, the choice of parameters significantly impacts performance across almost all sub-tasks. Therefore, depending on the specific task, selecting the appropriate values for $g$ and $h$ is crucial when using the OLP, as illustrated by the performance at high $\eta$ with non-optimal parameters in the bottom panel of Fig. 4.
IV.2 Considering a finite amount of resources
In the previous section, we analyzed the effect of making use of a weak measurement-based approach to QRC in the ideal scenario of infinite resources. Here we consider the case in which each expectation value extracted from the reservoir is computed with a finite number of shots $N_{s}$. To take this into account, it has been shown that it is enough to add a Gaussian noise dependent on $N_{s}$ and $g$ to the previously computed expectation values [mujal2023time]. In this context, it is important to consider the inherent inefficiency of the RSP compared to the OLP, as discussed in [mujal2023time]. This analysis takes into account the time resources required for executing each protocol. With the RSP, the time required to run the QRC scales quadratically with respect to the dimension of the time series, whereas for the OLP, this dependence is linear. Therefore, by fixing the number of shots available for the OLP, we can compute the required time and subsequently determine the number of shots available for the RSP within that same time frame. Approximately, we find that the number of shots available for the RSP is proportional to the number available for the OLP divided by $K$, the size of the dataset (see Appendix).
At this point, we can proceed to look at $P_{R}$, with the same logic as in Figure 2, but in a finite resources scenario. We now expect to see the combined effect of the back-action introduced by the OLP and of the higher efficiency of the OLP with respect to the RSP (Fig. 5). In this analysis, we fixed $N_{s}^{\text{OLP}}=1.5\cdot 10^{6}$. Again the $P_{R}$ is probed in $12\times 40$ points in the $g-h$ parameter space, and each point is the average over 50 different reservoir realizations.
Figure 5: Performance ratio analysis for the forward prediction task (top) and for the short-term memory task (bottom) in the scenario of finite resources. We fixed $N_{s}^{\text{OLP}}=1.5\cdot 10^{6}$ and computed $N_{s}^{\text{RSP}}$ consequently for both tasks ($\approx 1.5\cdot 10^{3}$ for the forward prediction task and $\approx 3\cdot 10^{3}$ for the short-term memory task).
In this case, the same clear message emerges from the analysis of the $P_{R}$ in both tasks: in the finite resources scenario defined above, using the OLP with properly tuned $g$ and $h$ leads to higher performance with respect to the RSP. This conclusion is conveyed by the red regions that appear in the plots for $h\geq 10^{-1}$ and $g\leq 1$. In these regions, the $P_{R}$ is substantially greater than $1$, meaning that the OLP is performing better than the best-performing RSP. This, as pointed out in the previous sections, is a combined effect of both the OLP being more efficient in terms of resources and of the positive effect that the back-action brings to the performance of the reservoir in the optimal regime for the measurement strength.
V Conclusions and Discussion
In this work, we analyzed the effects of weak measurements on the QRC algorithm through a comprehensive comparison between the OLP and the RSP. Our findings reveal that, contrary to previous assumptions, the unwanted back-action from weak measurements can positively influence the performance of QRC algorithms. By optimizing the performance ratio $P_{R}$, we highlighted that an informed choice of measurement strength and external fields in the OLP allows overperform the commonly used RSP. The study is performed on several sub-tasks derived from two main benchmark tasks, widely used in the literature. Furthermore, we show that an optimal parameter choice for the OLP can be crucial to obtaining good performances depending on specific sub-tasks.
As the field transitions to experimental implementations, our results highlight the necessity of identifying optimal protocols, especially in the finite resources scenario. In this context, we combine the previous knowledge on the better scaling of the OLP compared to the RSP with the ability to positively affect the reservoir dynamics by tuning the measurement strength. Our findings highlight the clear advantage of the OLP in this setting. This practical orientation is essential as experimental interest in QRC continues to grow [Yelin2024, Chen2020, Suzuki2022, spagnolo2022experimental, Pfeffer2022, Kubota2023, Yasuda2023, hu2024overcoming, Paternostro2024, PhysRevApplied.20.014051], providing a clear roadmap for researchers aiming to harness the capabilities of quantum systems for data processing tasks.
Moreover, our work addresses the growing concerns regarding the scalability and trainability of quantum machine learning models. As variational quantum algorithms face challenges related to these issues, QRC emerges as a compelling alternative. Not only by following this approach trainability issues such as barren plateaus [LaroccaBarrenPlateaus] are avoided. Using the optimized OLP we provide a pathway for conducting machine learning tasks efficiently at scale.
In conclusion, our analysis underscores the fundamental role of weak measurements in enhancing the performance of QRC through the OLP framework. By demonstrating that making use of weak measurements not only improves the scaling of the algorithm but can also improve its performance given an appropriate parameter selection, we lay the groundwork for future research.
Looking forward, we recognize the need for extended studies across a broader range of tasks, classical and quantum [Nokkala2023], and metrics, which may reveal even more generalized conclusions regarding the behavior of QRC algorithms under the OLP. In this direction, studying the effects of evolution time and coherence dynamics may provide deeper insights into the quantum processes underlying these systems. Our results strongly support the experimental realization of QRC algorithm with weak measurements through the OLP, particularly within emerging platforms that utilize neutral atoms and photons. As quantum computing evolves, these insights will drive both theoretical and practical applications in quantum machine learning.
Data and code availability statement
A tutorial version of our implementation of the QRC algorithm utilized in this work, along with the necessary data to partially reproduce some of the figures, is available at [git_tuto]. The complete dataset can be provided upon request.
Acknowledgements
We would like to express our gratitude to Rodrigo Martínez-Peña for their attentive reading and insightful comments on this manuscript. P.M. also thanks Rodrigo Martínez-Peña, Gian Luca Giorgi, Miguel C. Soriano, and Roberta Zambrini for their stimulating discussions at IFISC in Mallorca, which contributed significantly to the success of this project. G.F. would also like to thank Roberto Tricarico and Lisa Bombieri for the careful reading of the manuscript and the fruitful discussions. G.F. acknowledges support from a ”la Caixa” Foundation (ID 100010434) fellowship. The fellowship code is LCF/BQ/DI23/11990070. This project has received funding from MICIN, the European Union, NextGenerationEU (PRTR-C17.I1), the Government of Spain (Severo Ochoa CEX2019-000910-S and FUNQIP), the ERC AdG CERQUTE and the AXA Chair in Quantum Information Science, the ERC AdG NOQIA, MCIN/AEI (PGC2018-0910.13039/501100011033, CEX2019-000910-S/10.13039/501100011033, Plan National FIDEUA PID2019-106901GB-I00, Plan National STAMEENA PID2022-139099NB, I00, the “European Union NextGenerationEU/PRTR" (PRTR-C17.I1), FPI, QUANTERA MAQS PCI2019-111828-2, QUANTERA DYNAMITE PCI2022-132919, QuantERA II Programme co-funded by European Union’s Horizon 2020 program under Grant Agreement No 101017733; Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program, AGAUR Grant No. 2021 SGR 01452, QuantumCAT  U16-011424, co-funded by ERDF Operational Program of Catalonia 2014-2020); Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024); Funded by the European Union.
Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, European Commission, European Climate, Infrastructure and Environment Executive Agency (CINEA), or any other granting authority. Neither the European Union nor any granting authority can be held responsible for them (HORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690, EU Horizon 2020 FET-OPEN OPTOlogic, Grant No 899794), EU Horizon Europe Program (This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101080086 NeQSTGrant Agreement 101080086 — NeQST); ICFO Internal “QuantumGaudi” project; European Union’s Horizon 2020 program under the Marie Sklodowska-Curie grant agreement No 847648; “La Caixa” Junior Leaders fellowships, La Caixa” Foundation (ID 100010434): CF/BQ/PR23/11980043.
Appendix
Details on the finite resource scenario
Here we consider the case in which each expectation value extracted from the reservoir is computed with a finite number of shots $N_{s}$. To take this into account, it has been shown that it is enough to add a Gaussian noise dependent on $N_{s}$ and $g$ to the previously computed expectation values [mujal2023time]. Namely, for one and two spins observables the following statistical uncertainties are considered:
$\overline{s}_{\hat{\sigma}}=\sqrt{\frac{g^{2}+1}{g^{2}N_{s}}},\quad\overline{s}_{\hat{\sigma}\otimes\hat{\sigma}}=\sqrt{\frac{g^{4}+2g^{2}+1}{g^{4}N_{s}}}.$ (7)
Within this scenario, it is also important to consider the inherent inefficiency of the RP compared to the OLP as discussed in [mujal2023time]. This analysis builds upon taking into account the time resources that each protocol requires to be executed. Let’s start by assuming that the time required to reset the reservoir and to take a measurement is negligible. Hence, we can define the experimental time required to run the QRC algorithm with the RSP on a dataset of $K$ elements as:
$T_{\text{exp}}^{\text{RSP}}=3N_{s}^{\text{RSP}}\left(K^{\prime}\tau_{wo}+\frac{1}{2}(K^{\prime}+1)K^{\prime}\Delta t\right),$ (8)
where $K^{\prime}$ is equal to $K$ minus the number of samples used for the washout stage $K_{wo}$ and $\tau_{wo}=K_{wo}\Delta t$ is the time required for the washout stage. On the other hand, in the case of using the OLP we get a linear dependence on $K^{\prime}$ rather than a quadratic:
$T_{\text{exp}}^{\text{OLP}}=3N_{s}^{\text{OLP}}\left(K\Delta t\right).$ (9)
In both (8), (9) the factor $3$ is since we cannot measure simultaneously in all the directions $x$, $y$ and $z$.
Then, if we fix $N_{s}$ for the OLP, we expect to have a smaller amount of resources for the RSP given the assumption that we have the same experimental time at our disposal for both protocols. To see this, let’s formally set $T=T_{\text{exp}}^{\text{OLP}}=T_{\text{exp}}^{\text{RSP}}$ as the available resource in time and fix the number of shots for the OLP. The available shots for the RSP result to be:
$N_{s}^{\text{RSP}}=N_{s}^{\text{OLP}}\frac{2K}{K^{\prime 2}+K^{\prime}(2K_{wo}+1)}.$ (10)